# A novel characterization of structures in smooth regression curves: from a viewpoint of persistent homology

### v3

@@ 8.4 Some Auxiliary Results
For the random variable $Z_{i}$ defined in the Equation 8.3, for large enough $n$ such that $h\equiv h(n)$ is sufficiently small, then we have $\mathbb{E}[Z_{i}]\lessapprox-\epsilon\delta p_{min}2h$.
@@
[…]
$=p_{max}\left(M-m(x)\right)\left[K_{h}\left(-\tau h\right)-K_{h}\left(\tau h\right)\right]-\delta p_{min}\left[m(x+h)-m(x-h)\right]-\epsilon\delta p_{min}2h$
@@
$\approx-\epsilon\delta p_{min}2h,$
where the last line follows from assuming that for small enough h, we $h,\text{ we have }m(x+h)\approx m(x-h)$. $\Box$
Acknowledgment : The authors thank the editor, an anonymous associate editor, and two anonymous reviewers for their excellent comments and constructive criticisms, which significantly improve the article. Moreover, Subhra Sankar Dhar gratefully acknowledges his core research grant (Grant Number: CRG/2022/001489), the Government of India.

### v4

@@ 8.4 Some Auxiliary Results
For the random variable $Z_{i}$ defined in the Equation 8.3, we have:
$\mathbb{E}[Z_{i}]\leq-2h\delta p_{min}(\gamma_{1}+\epsilon)$
@@
$=\int_{x-h}^{x+h}K_{1,h}(x-\xi)\left(m(\xi)-m(x)\right)p(\xi)d\xi-\int_{x-h}^{x+h}\left(m_{1}\left(\xi\right)+\epsilon\right)K_{h}(x-\xi)p(\xi)d\xi$
$\leq p_{max}\left(M-m(x)\right)\int_{x-h}^{x+h}K_{1,h}(x-\xi)d\xi-\delta p_{min}\int_{x-h}^{x+h}m_{1}\left(\xi\right)d\xi-\epsilon\delta p_{min}2h$
$=p_{max}\left(M-m(x)\right)\left[K_{h}\left(-h\right)-K_{h}\left(h\right)\right]-\delta p_{min}\left[m(x+h)-m(x-h)\right]-\epsilon\delta p_{min}2h$
@@
$\leq-2h\delta p_{min}(\gamma_{1}+\epsilon),$
where the last line follows from the application of the mean value theorem for $m$ on the interval (xh,x+h)(x-h,x+h), and then subsequently using Assumption (B.5). $\Box$
8.5 Comparison and Implementation of Algorithm 1
In this subsection, we implement Algorithm 1 on the simulated data sets used in Section 5 to investigate whether the estimated derivative values are reasonable or not. Note that, for our purpose, some reasonable bounds of the unknown population quantities are sufficient rather than the exact values. We compare the derivative values estimated from Algorithm 1 with the true values and values estimated from the ratio of first-order differences of Y(response) and X(regressor). We denote $\hat{Y}_{1}^{T},\hat{Y}_{1}^{A}\text{ and }\hat{Y}_{1}^{N}$ as the sets that contain true derivative values, derivative values estimated from Algorithm 1, and derivative values estimated from the ratio of first-order differences of Y and X, respectively.
We simulate $n=1000$ observations from the functions used in Section 5.1, Section 5.2, and Section 5.3 with the same error distribution. We denote simulated data sets, for $n=1000$, used in Section 5.1, Section 5.2, and Section 5.3 as $D_{1}$, $D_{2}\text{, and }D_{3}$, respectively. We provide summary statistics of the estimated derivative values and plot the estimated values for visual comparison with the true values. We denote the first and third quartiles by $\boldsymbol{Q_{1}}$ and $\boldsymbol{Q_{3}}$, respectively, in the summary tables. The following are the results for the data sets $D_{1},D_{2}\text{ and }D_{3}$.
Results for $\boldsymbol{D_{1}}$: It is evident from Table 2 and Figure 14 that Algorithm 1 provides good estimates except for the lower extreme of the true values. On the other hand, the values in $\hat{Y}_{1}^{N}$ are too noisy.
Table 2: Summary statistics for $D_{1}$
Values Min. $\boldsymbol{Q_{1}}$ Median Mean $\boldsymbol{Q_{3}}$ Max.
$\hat{Y}_{1}^{T}$ 0.3693 0.6018 0.9914 1.1620 1.6033 2.7108
$\hat{Y}_{1}^{A}$ 0.0945 0.7538 1.0169 1.0621 1.3013 2.1057
$\hat{Y}_{1}^{N}$ -730.0287 0.7685 1.0332 0.3342 1.3647 5.6000
Figure 14: Plots for $D_{1}$.
Results for $\boldsymbol{D_{2}}$: It is evident from Table 3 and Figure 15 that Algorithm 1 provides good estimates except for the upper extreme of the true values. On the other hand, the values in $\hat{Y}_{1}^{N}$ are too noisy in this case as well.
Table 3: Summary statistics for $D_{2}$
Values Min. $\boldsymbol{Q_{1}}$ Median Mean $\boldsymbol{Q_{3}}$ Max.
$\hat{Y}_{1}^{T}$ -1.9925 -1.0156 -0.0173 -0.3357 0.3503 0.4593
$\hat{Y}_{1}^{A}$ -1.8059 -0.5828 -0.1350 -0.2103 0.2641 1.4762
$\hat{Y}_{1}^{N}$ -731.5696 -0.5851 -0.1371 -1.0017 0.2696 3.3662
Figure 15: Plots for $D_{2}$.
Results for $D_{3}$: It is evident from Table 4 and Figure 16 that Algorithm 1 provides good estimates of true values around 0. However, the values at both extremes of the true values are not estimated well. Here also the values in $\hat{Y}_{1}^{N}$ are too noisy.
Table 4: Summary statistics for $D_{3}$
Values Min. $\boldsymbol{Q_{1}}$ Median Mean $\boldsymbol{Q_{3}}$ Max.
$\hat{Y}_{1}^{T}$ -12.10 -1.602 0.0007 0.2759 2.305 12.10
$\hat{Y}_{1}^{A}$ -3.7409 -0.5490 -0.0002 -0.0060 0.5035 3.8894
$\hat{Y}_{1}^{N}$ -741.4168 -0.9544 -0.0012 -0.8033 0.7530 11.2697
Figure 16: Plots for $D_{3}$.
Acknowledgments
The authors thank the editors, an anonymous associate editor and two anonymous reviewers for their excellent comments and constructive criticisms, which improved the article significantly. Subhra Sankar Dhar gratefully acknowledges his core research grant (Grant Number: CRG/2022/001489), from the Government of India.